# Data Poisoning for In-context Learning

### v2

@@ 3.3 Attacking Algorithms
The perturbation $\delta$ is a crucial component of the ICLPoison framework, designed to be imperceptible to humans while effective in manipulating the performance. These are also common requirements for NLP attacks (Ebrahimi et al., 2017; Jin et al., 2020; Xu et al., 2023; Li et al., 2018). On the other hand, the discrete nature of the objective in Eq.3 poses additional challenges in the optimization. To address these challenges as well as showcase the versatility of our framework, we introduce three representative perturbations: synonym replacement, character replacement, and adversarial suffix. These methods demonstrate the adaptability of ICLPoison across different levels of text manipulation: Synonym replacement evaluates the word-level vulnerability of ICL and subtly changes the semantics; character replacement involves minimal but precise alterations, such as changing individual letters, making it less noticeable to human reviewers (see examples in Appendix A.3); and adversarial suffix test token-level vulnerabilities in ICL, accommodating different models by adapting to their unique tokenization methods. In general, these perturbations allow for a comprehensive evaluation of the vulnerability in ICL. The optimization of these perturbations is efficiently managed through a greedy search method, proven effective in similar contexts (Lei et al., 2019; Bao et al., 2022).
Synonym Replacement. This approach involves substituting words in a text with their synonyms, aiming at preserving the semantic meaning and grammatical structure (Jin et al., 2020; Xu et al., 2023). Therefore, it serves as an appropriate example of $\delta$ in Section 3.2. Within our framework, we limit the number of word replacements (denoted as $k$) to maintain the perturbation’s imperceptibility. For a text composed of a sequence of $n$ words $x=[w_{1},...,w_{n}]$, $\delta(x)$ is defined as $[s_{1},...,s_{n}]$, where $s_{i}$ is either a synonym of $w_{i}$ (if selected for replacement) or remains as $w_{i}$ (if not replaced). Our objective is to identify the optimal replacements that maximize the objective in Eq.3. This brings natural challenges: which words are to be replaced and what synonyms are suitable. Following a strategy similar to (Jin et al., 2020), we adopt a two-step optimization process. Firstly, we calculate an importance score for each word, selecting the top-$k$ words with the highest scores for replacement. The importance score for word $w_{i}$ is computed as the distortion (in Eq.2) before and after deleting $w_{i}$, expressed as:
@@
where $x_{\text{\textbackslash}w_{i}}$ denotes the text after the removal of $w_{i}$. In the second step, we use a greedy search method. This method involves iteratively finding the best replacement for each selected word, one at a time, while keeping the other words fixed (Yang et al., 2020; Lei et al., 2019). In each iteration, we focus on one of the selected words, replacing it with its most appropriate synonym. We retrieve synonyms using GloVe word embeddings (Pennington et al., 2014), selecting those with embeddings most similar (based on cosine similarity) to the original word. After identifying potential synonyms, each synonym is temporarily substituted into the text, and the loss function in Eq.3 is evaluated. The synonym that results in the highest loss is then chosen as the final replacement. We then proceed to the next word, repeating this replacement process until traversing all selected words. We present the detailed algorithm in Algorithm. 1.
@@
Character Replacement. This method is similar with the synonym replacement approach, but focuses on replacing individual characters instead of whole words (Ebrahimi et al., 2017; Lei et al., 2019; Xu et al., 2023). When changing only a few letters, this method can be less detectable to humans and maintain the word’s pronunciation and basic structures (Ebrahimi et al., 2017). We limit the number of character replacements to $k$ to maintain the subtlety of the perturbations. The optimization process for character replacement also follows a two-step method. Firstly, we calculate the importance score for each character, similar to the approach described in Eq.4. The primary difference is that we consider the removal of individual characters, not words. The top-$k$ characters with the highest importance scores are earmarked for replacement. Secondly, we employ the same greedy search strategy used in synonym replacement to identify the optimal replacements for the selected characters. Note that our character set encompasses uppercase and lowercase letters, digits, punctuation marks, and whitespace, in line with the sets in (Kim et al., 2016; Ebrahimi et al., 2017). The detailed algorithm and its implementation are shown in Algorithm 2 in Appendix  A.1.
@@
Adversarial Suffix. The concept of an adversarial suffix, referred to as adding additional tokens at the end of the original text, has shown considerable effectiveness in misleading LLMs (Zou et al., 2023). Thus, in addition to synonym and character replacement, we also adapt this perturbation to evaluate the token-level vulnerability of the ICL process. To ensure imperceptible to humans, we restrict the number of additional tokens when adapting to our framework. For a given text $x$ that can be tokenized into a sequence of tokens $x=[t_{1},...,t_{n}]$, we define $\delta(x)$ as $[t_{1},...,t_{n},t^{\prime}_{1},...,t^{\prime}_{k}]$ where $t^{\prime}_{1},...,t^{\prime}_{k}$ are adversarial suffices. Our goal is to identify the optimal suffixes that maximize the objective in Eq 3. We also employ a greedy search approach, involving iteratively selecting each suffix token from $t^{\prime}_{1}$ to $t^{\prime}_{k}$ one by one which results in the maximum increase in the loss. The detailed implementation and optimization process for this approach is elaborated in Algorithm 3 in Appendix  A.1.
4 Experiment
@@
We conduct extensive experiments to validate the effectiveness of the proposed framework ICLPoison, particularly with three perturbations introduced in Section 3.3.
@@ 4.1 Experiments setting
Datasets. We conduct experiments on different types of classification tasks. Stanford Sentiment Treebank (SST2) dataset from the GLUE benchmark (Wang et al., 2019) is a sentiment analysis dataset consisting of sentences from movie reviews and human annotations of their sentiment in 2 classes (positive/negative); Corpus of Linguistic Acceptability (Cola) dataset from GLUE is a linguistic analysis dataset consisting of English acceptability judgments collected from linguistic books, labeled with “acceptable” or “unacceptable”; Emo dataset (Wang et al., 2023) focuses on emotion classification consisting of Twitter messages labeled in 4 classes; AG’s new (AG) corpus (Zhang et al., 2015) is a topic classification dataset gathered from news sources and labeled in 4 classes (World, Sports, Business, and Science/Technology); and Poem Sentiment (Poem) (Sheng & Uthus, 2020) is a sentiment analysis dataset of poem verses from Project Gutenberg, classified into 3 classes.
Models. In this work, we consider both open-source models including Llama2-7B (Touvron et al., 2023), Pythia (2.8B, 6.9B) (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023), GPT-J-6B (Wang & Komatsuzaki, 2021), MPT-7B (Team, 2023), and API-only models GPT-3.5 and GPT-4 (Brown et al., 2020).
@@
Metrics. Our main focus is on the ICL accuracy. For every dataset, we generate perturbations for examples in the training data and randomly select examples from it to conduct ICL prediction for every sample in the test data. We repeat for 5 runs and report the average ICL accuracy. We also include perplexity scores, which are defined as the average negative log-likelihood of each of the tokens appearing, showing whether the perturbations are imperceptible or not.
@@
4.2 Main results
In this subsection, we apply our framework to various LLMs and datasets to evaluate the vulnerability of ICL.
Attacking open-source models. Our study initially examines open-source models. For each, we craft poisoned samples utilizing the model’s own architecture and subsequently assess the ICL accuracy. Note that we will discuss the transferability i.e. examining the performance of poisoned samples crafted using Llama2-7B when tested on different models in Section 4.3. We begin by altering the entire training set to create these samples. Further studies involving only a subset of poisoned data are presented in Section 4.4. Partial results are shown in Table 1 and full results can be found in Table 5 in Appendix A.2. In the table, a lower accuracy indicates a stronger poisoning effect and the lowest performance is highlighted. It is obvious that ICL performs well on clean data, especially for the SST2 dataset and Llama2-7B model, achieving more than 88% accuracy. While random labels do interrupt the ICL process, causing a decrease in performance, LLMs are still capable of learning effectively from the demonstrations with the dropping of ICL accuracy less than 7%, which is aligned with observations by (Min et al., 2022). In contrast, our ICLPoison framework significantly reduces ICL accuracy, achieving drops to below 10% for some models and datasets such as GPT-J-6B with the Emo dataset. Notably, ICLPoison reveals that ICL is vulnerable to different levels of data poisoning attacks, ranging from subtle character-level changes to word-level manipulations. Among all three variants of ICLPoison, synonym replacement and adversarial suffix perturbations exhibit more severe threats than character replacement in reducing ICL accuracy, likely because they introduce more changes within the same perturbation budget. We present some poisoned examples from three methods in Appendix A.3 for human evaluation. We observe that the poisoning effect differs across models and datasets. For instance, Llama2-7B shows heightened sensitivity to poisoning on the Emo dataset, whereas Falcon-7B is particularly vulnerable to attacks on the Poem Sentiment dataset. These variations may stem from differences in model architectures and the inherent complexity of the datasets.
@@ 4.2 Main results
4.3 Transferbility
In this subsection, we consider a practical scenario when the attacker has no access to the victim LLM, and leverages a surrogate model to generate poisoned samples. Specifically, we adopt the Llama2-7B model in this study. This setup allows us to evaluate the transferability of our poisoning approach across different models, including black-box models such as GPT-3.5 and GPT-4. The initial results for the transferability study are presented in Figure 2, in which we focus on the SST2 dataset. For a more comprehensive analysis covering all five datasets used in our study, we direct readers to Appendix A.2.
@@ 4.3 Transferbility
From Figure 2, we observe a notable trend: while the efficacy of the poisoning attack diminishes when moving from the surrogate (Llama2-7B) to the victim models, the impact remains significant. The poisoned examples generated by our ICLPoison framework – in all three of its variants – substantially jeopardize the performance of ICL, leading to over a 30% decrease in accuracy for open-source models. Furthermore, our analysis sheds light on the different transferability of various perturbations. Synonym replacement and adversarial suffixes, in particular, demonstrate a stronger poisoning effect compared to character replacement. This disparity could be attributed to the more pronounced influence these methods exert on the surrogate model. We also note the influence of model size on susceptibility to poisoning. Models smaller in size than the surrogate, such as Pythia-2.8B and GPT-J-6B, appear to be more vulnerable to these poisoning examples, while larger models exhibit a degree of resistance. This pattern suggests that the effectiveness of our approach may be modulated by the size and complexity of the target model, offering valuable insights for future research in trustworthy language models.
4.4 Partial poisoning
To consider practical scenarios where attackers can only alter a part of the data, our experiments involve perturbing a random subset of the training dataset. We allocate varying proportions (10%, 20%, 50%, and 100%) of the training data as accessible for manipulation in the Llama2-7B model and GLUE-SST2 dataset. The results, shown in Figure 3, reveal that a lower mixing rate reduces the overall poisoning effect. Nonetheless, even at a 10% rate, we observe a significant decrease in ICL performance by over 7%, and a 15% decrease at 20%, underscoring the efficacy of our ICLPoison framework.
@@ 4.4 Partial poisoning
Figure 3: Results for partially poisoned ICL. The X-axis represents the ratio of poisoned data while Y-axis represents the ICL accuracy.
These findings highlight the vulnerability of ICL to subtle data poisoning attacks, where even a limited number of malicious inputs can significantly disrupt the ICL process. Additionally, the results show that perturbation strategies like synonym replacement and adversarial suffix, which introduce more pronounced textual changes within a fixed poisoning budget, more severely affect ICL performance compared to character replacement. This indicates that broader, coarse-grained perturbations are generally more disruptive than finer, more subtle ones.
@@
4.5 Potential defenses
To evaluate the robustness of our framework, we applied two representative defenses from (Jain et al., 2023): detection-based defense perplexity filter and preprocessing defense paraphrasing.
Perplexity filter. The perplexity score of a text is referred to as the average negative log-likelihood of each of the tokens appearing. The perturbation added to the original text may cause grammar mistakes, logic problems, and a reduction in fluency. This will increase the perplexity, and thus the corresponding text could be more detectable (Jain et al., 2023). We report the perplexity scores of generated poisoned data, and a higher value indicates a higher probability of being detected and lower robustness. Table 2 shows the perplexity scores for poisoned data across various models and datasets, calculated as described in Section 4.1 of (Jain et al., 2023). Among the perturbations, synonym replacement yields relatively lower perplexity most close to the clean perplexity, whereas adversarial suffixes generally result in the highest scores. This suggests that synonym replacement is less detectable than the adversarial suffix.
Paraphrasing. Paraphrasing is a typical preprocessing method that utilizes a language model to rewrite the input text. This process aims to maintain original meanings while removing adversarial perturbations, making it an effective defensive strategy. In this part, we leverage the cutting-edge GPT-4 model to paraphrase the poisoned data and evaluate ICL performance on paraphrased inputs. As shown in Table 3, paraphrasing successfully neutralizes adversarial suffixes while largely preserving the impact of synonym replacements. This outcome is anticipated since adversarial suffixes typically introduce irrelevant content, whereas synonym replacements maintain the text’s semantic integrity.
Table 3: The poisoned data is paraphrased by GPT-4, and ICL accuracy is reported. Original results are included in brackets.
@@ 4.5 Potential defenses
The above findings make it evident that the implemented defenses can enhance the robustness of ICL against data poisoning to some extent. However, the degree of improvement is contingent upon the type of perturbation. For instance, token-level perturbations, such as adversarial suffixes, are more readily addressed by these defenses, whereas word-level perturbations like synonym replacements pose a greater threat due to their subtlety and ability to alter semantic meaning. Character-level perturbations fall between these two in terms of detectability and potential impact, with their severity likely to increase with a higher attack budget. Our results highlight the necessity to develop a robust ICL.

### v3

@@ 4.3 Attacking Algorithms
To design the perturbation $\delta$, as required in common NLP attacks (Ebrahimi et al., 2017; Jin et al., 2020; Xu et al., 2023; Li et al., 2018), $\delta$ is supposed to be imperceptible to humans while effective in manipulating the performance. In addition to the above requirements, an additional challenge to consider in the optimization is the discrete nature of the objective in Eq.2. To address these requirements and challenges as well as showcase the versatility of our method, we introduce three representative perturbations: synonym replacement, character replacement, and adversarial suffix. These methods demonstrate the adaptability of ICLPoison across different levels of text manipulation: Synonym replacement evaluates the word-level vulnerability of ICL and subtly changes the semantics; character replacement involves minimal but precise alterations, making it less noticeable to human reviewers (see examples in Appendix A.4); and adversarial suffix test token-level vulnerabilities in ICL. The optimization of these perturbations is managed through a greedy search method, proven effective in similar contexts (Lei et al., 2019; Bao et al., 2022).
Synonym Replacement. This approach involves substituting words in a text with their synonyms, aiming at preserving the semantic meaning and grammatical structure (Jin et al., 2020; Xu et al., 2023). Within our method, we limit the number of word replacements (denoted as $k$) to maintain the perturbation’s imperceptibility. For a text composed of a sequence of $n$ words $x=[w_{1},...,w_{n}]$, $\delta(x)$ is defined as $[s_{1},...,s_{n}]$, where $s_{i}$ is either a synonym of $w_{i}$ (if selected for replacement) or remains as $w_{i}$ (if not replaced). To identify which words are to be replaced and which synonyms are suitable, we follow a strategy similar to (Jin et al., 2020). We adopt a two-step optimization process. First, we calculate an importance score for each word, selecting the top-$k$ words with the highest scores for replacement. The importance score for word $w_{i}$ is computed as the distortion (in Eq.1) before and after deleting $w_{i}$, expressed as:
@@
where $x_{\text{\textbackslash}w_{i}}$ denotes the text after the removal of $w_{i}$. In the second step, we use a greedy search method that iteratively replaces each selected word while keeping others fixed (Yang et al., 2020; Lei et al., 2019). For each word, we find synonyms using GloVe embeddings (Pennington et al., 2014), selecting those with the highest cosine similarity to the original word. Each synonym is temporarily substituted into the text, and the loss function in Eq.2 is evaluated. The synonym that maximizes the loss is chosen as the final replacement. This process is repeated for all selected words. We present the detailed algorithm in Algorithm. 1.
@@
Character Replacement. This method is similar to the synonym replacement approach but focuses on replacing individual characters instead of whole words (Ebrahimi et al., 2017; Xu et al., 2023). When changing only a few letters, this method can be less detectable to humans and maintain the word’s pronunciation and basic structures (Ebrahimi et al., 2017). We limit character replacements to $k$ to keep perturbations subtle. The optimization process involves two steps: first, we calculate an importance score for each character, similar to Eq.3 but focusing on individual character removal rather than words. The top-$k$ characters with the highest scores are selected for replacement. Second, we use a greedy search strategy, similar to synonym replacement, to find the best replacements for these characters. Note that our character set encompasses uppercase and lowercase letters, digits, punctuation marks, and whitespace, in line with the sets in (Kim et al., 2016; Ebrahimi et al., 2017). The detailed algorithm and its implementation are shown in Algorithm 2 in Appendix  A.1.
@@
Adversarial Suffix. The concept of an adversarial suffix, referred to as adding additional tokens at the end of the original text, has shown considerable effectiveness in misleading LLMs (Zou et al., 2023). Thus, in addition to synonym and character replacement, we also adapt this perturbation to evaluate the token-level vulnerability of the ICL process. To ensure imperceptible to humans, we restrict the number of additional tokens when adapting to our method. For a given text $x$ that can be tokenized into a sequence of tokens $x=[t_{1},...,t_{n}]$, we define $\delta(x)$ as $[t_{1},...,t_{n},t^{\prime}_{1},...,t^{\prime}_{k}]$ where $t^{\prime}_{1},...,t^{\prime}_{k}$ are adversarial suffices. Our goal is to identify the optimal suffixes that maximize the objective in Eq 2. We also employ a greedy search approach and iteratively select each suffix token from $t^{\prime}_{1}$ to $t^{\prime}_{k}$ one by one which results in the maximum increase in the loss. The detailed implementation and optimization process for this approach is in Algorithm 3 in Appendix A.1.
5 Experiments
@@
We conduct extensive experiments to validate the effectiveness of the proposed method ICLPoison, particularly with three perturbations introduced in Section 4.3.
@@ 5.1 Experiments setting
Datasets. We conduct experiments on different types of classification tasks. SST2 (2 classes) and Cola (2 classes) are from GLUE dataset (Wang et al., 2019); Emo (Wang et al., 2023) (4 classes) is an emotion classification dataset; AG’s new (4 classes) (Zhang et al., 2015) is a topic classification dataset; Poem (3 classes) (Sheng and Uthus, 2020) is a sentiment analysis dataset of poem. Details of these datasets are presented in Appendix A.2.
Models. We use open-source models including Llama2-7B (Touvron et al., 2023), Pythia (2.8B, 6.9B) (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) [^2] Two additional models GPT-J-6B (Wang and Komatsuzaki, 2021), MPT-7B (Team, 2023) in Appendix A.3, and API-only models GPT-3.5 and GPT-4.
@@
Metrics. Our main focus is on ICL accuracy. For every dataset, we generate perturbations for examples in the training data and randomly select examples from it to conduct ICL prediction for every sample in the test data. We repeat for 5 runs and report the average ICL accuracy. We also include perplexity scores, which are defined as the average negative log-likelihood of each of the tokens appearing, showing whether the perturbations are imperceptible or not.
@@
5.2 Effectiveness of ICLPoison
@@
In this subsection, we first validate the effectiveness of the idea of corrupting the hidden states in ICLPoison across various LLMs and tasks. To eliminate the potential impact on random sampling of poisoned/clean examples, we consider the extreme case of 100% poisoning rate [^4] This is an extreme case but still practical. For instance, the victim downloads a whole poisoned data set generated by the attacker for ICL prediction.. A more practical attacking scenario with small poisoning rates is discussed in Section 5.3.
Attacking open-source models. We first examine open-source models. For each, we craft poisoned samples utilizing the model’s own architecture and assess the ICL accuracy. Partial results are shown in Table 1 and full results can be found in Table 7 in Appendix A.3. In the table, a lower accuracy indicates a stronger poisoning effect, and the lowest performance is highlighted. One can see that ICL performs well on clean data, especially for the SST2 dataset and Llama2-7B model, achieving more than 88% accuracy. Besides, the random labeling can only decrease the accuracy marginally, less than 7%, which is aligned with observations by (Min et al., 2022). In contrast, our ICLPoison method significantly reduces ICL accuracy, achieving drops to below 10% for some models and datasets such as Falcon-7B with the Emo dataset. Notably, the effectiveness of ICLPoison implies that ICL is vulnerable to data poisoning attacks that corrupt the hidden states.
In addition, among the three variants of ICLPoison, synonym replacement and adversarial suffix perturbations pose more severe threats to ICL accuracy compared to character replacement. This disparity may arise because, within the same perturbation budget, different types of perturbations induce varying degrees of change in the hidden states. Character changes typically do not alter the semantic content as significantly as synonym replacements or adversarial suffixes, which can introduce more substantial shifts in the hidden states and disrupt the model’s prediction more effectively. We present some poisoned examples from three methods in Appendix A.4 for human evaluation.
@@
Transferbility. We adopt the Llama2-7B model as the surrogate model and test ICL performance on other models to evaluate the transferability of our poisoning approach across different models, including black-box models. This evaluation validates the transferability of proposed ICLPoison and reveals factors influencing the transferability such as perturbation type and model size.
@@
Initial results, presented in Figure 1, focus on the SST2 dataset, with a comprehensive analysis of all five datasets available in Appendix A.3. The results indicate that while the effectiveness of the poisoning attack diminishes when moving from the surrogate (Llama2-7B) to other models, the impact remains significant. Our ICLPoison method—across its three variants—leads to over a 30% decrease in accuracy for open-source models. This underscores the substantial threat posed by these attacks. Our analysis also reveals differences in how various perturbations transfer across models. Synonym replacement and adversarial suffixes demonstrate a stronger poisoning effect compared to character replacement, likely because they introduce more disruption to the model’s hidden states. Furthermore, smaller models, such as Pythia-2.8B and GPT-J-6B, are more susceptible to these poisoning examples, whereas larger models exhibit some resistance. This suggests that the effectiveness of our approach can be affected by the size and complexity of the target model.
@@
Figure 2: Results for practical ICLPoison. The X-axis represents the poison rate while Y-axis represents the ICL accuracy.
The above experimental results confirm that distortions in hidden states can significantly compromise ICL predictions. Synonym replacements and adversarial suffixes cause greater changes in hidden states, leading to stronger poisoning effects. These distortions are also transferable, with smaller models being more affected. These insights underscore the critical link between hidden states and ICL predictions, emphasizing the effectiveness of data poisoning via manipulating hidden states.
5.3 Practical Attacks
To align with the threat model in Section 4.1, we consider practical attacking scenarios when attackers only insert some poisoned data into datasets. This usually happens when the victim collects data from various sources and mixes the poisoned data uploaded by the attacker with other clean data (Steinhardt et al., 2017; Carlini et al., 2024). In our experiments, we mix the poisoned data with clean training data, and the poisoning rate varies (10%, 20%, 50%, and 100%). We test with the Llama2-7B model and GLUE-SST2 dataset. The results, shown in Figure 2, reveal that ICLPoison is still a threat even when only a small set of poisoned data is injected into the clean prompt set. At a small poisoning rate such as 10% and 20%, we can observe a significant decrease in ICL performance by over 10% and 19% respectively, underscoring the efficacy of our proposed ICLPoison. These findings highlight the vulnerability of ICL to subtle data poisoning attacks, where even a limited number of malicious inputs can significantly disrupt the ICL process. Additionally, similar to the observations in Table 1, the results show that perturbation strategies like synonym replacement and adversarial suffix, which introduce more pronounced textual changes within a fixed poisoning budget, more severely affect ICL performance compared to character replacement.
5.4 Potential Defenses
Since the attackers are not aware of how victims leverage the poisoned data and some naive defenses may be applied, it is necessary to evaluate the robustness of an attack method. To examine the robustness of ICLPoison, we evaluate three representative defenses: perplexity filtering, paraphrasing (Jain et al., 2023), and grammar check. We also take 100% poisoning rate as the illustration and apply each defense.
@@ 5.4 Potential Defenses
Perplexity filtering. The perplexity score measures the average negative log-likelihood of tokens in a text. Perturbations to the original text can increase this score, making them potentially detectable (Jain et al., 2023). We present the perplexity scores of generated poisoned data using the Llama2-7B model; higher scores suggest a greater chance of detection and defense. Table 2 displays these scores for poisoned data across various models and datasets, following the calculation method in Section 4.1 of (Jain et al., 2023). Among perturbations, synonym replacement results in perplexity scores closest to clean data, while adversarial suffixes yield the highest scores.
Paraphrasing. Paraphrasing is a preprocessing technique that uses a language model to rewrite text, aiming to preserve original meanings while removing adversarial perturbations. We use GPT-4 to paraphrase poisoned data and assess ICL performance on these inputs with the Llama2-7B model. Table 3 shows that paraphrasing effectively neutralizes adversarial suffixes but largely retains the effects of synonym replacements. To explain this, adversarial suffixes add irrelevant content, while synonym replacements preserve semantic integrity [^5] Paraphrased text examples are provided in Appendix A.5..
Table 3: ICL accuracy when paraphrased by GPT-4. Original results are included in brackets.
@@
These findings indicate that defenses enhance ICL’s robustness against data poisoning, but their effectiveness varies by perturbation type. Token-level perturbations like adversarial suffixes are easier to counter, whereas word-level synonym replacements are more challenging due to subtle semantic changes. Character-level perturbations are moderately detectable but become more severe with larger attack budgets. These results highlight the need for more robust ICL defenses.
Grammar check defenses. We evaluate the performance of ICLPoison against typo-corrector (Pruthi et al., 2019) and popular grammar-checking tools–Grammarly and Microsoft Editor. We conduct experiments on SST2 dataset and Llama2-7B model with all three attacking strategies. For typo-corrector, we report the average ICL accuracy; for Grammarly, we report the ratio of correctness errors (CER) over the total number of words; for Microsoft Editor, we report the ratio of spelling errors (SER) and the ratio of grammar errors (GER) over the total number of words. As shown in Table 4, the proposed attack can bypass the grammar/typo detection method, especially the synonym replacement strategy. This highlights the stealthiness of ICLPoison.
Table 4: Attacking performance against detection defenses on model Llama2-7B and SST2 dataset.
Defenses Clean Synonym Adv Suffix Character
Grammarly
(CER ($\downarrow$))
4.89 5.04 11.97 10.31
Microsoft editor
(SER/GER ($\downarrow$))
2.72/0.62 2.68/0.65 10.08/0.63 7.25/0.68
Typo-corrector
(ICL accuracy ($\downarrow$))
90.2 23.7 41.4 35.8